OpenStudy (anonymous):

About 1/4 of all adults in the United States have type O+ blood. If four randomly selected adults donate blood, find the probability of each of the following events. (Enter your answers as fractions.) (a) All four are type O+. (b) None of them is type O+. (c) Two out of the four are type O+.

OpenStudy (anonymous):

let A be the event that adults have blood type O a)P(A)=1 B)P(A')=1-P(A)=0 c)P(2/4)=1-0.5=0.5